OpenStudy (anonymous):

solve the following differential equation- (p^2)+2pycotx=(y^2) where p=dy/dx

OpenStudy (anonymous):

chapter ka nam hai diffrential eqn of first order and higher degree

OpenStudy (usukidoll):

method of undetermined coeficcents? !

OpenStudy (experimentx):

find the value of 'p' in that quadratic ... see what you get. or maybe you can reduce it to http://en.wikipedia.org/wiki/Clairaut's_equation

OpenStudy (anonymous):

plz colud u solve it widout the help of clairauts eqn

OpenStudy (experimentx):

according to MMA solution is \[ \left\{\left\{y(x)\to c_1 \csc ^2\left(\frac{x}{2}\right)\right\},\left\{y(x)\to c_1 \sec ^2\left(\frac{x}{2}\right)\right\}\right\} \]

OpenStudy (usukidoll):

oops prof didn't teach me that.. sorry ^^

OpenStudy (experimentx):

method of undetermined coefficients is for linear equation. this is non linear equation.

OpenStudy (anonymous):

ok thnk u! first i have to thorugh the topic.

OpenStudy (experimentx):

how do you solve quadratic equation? .... use that for p^2 + a p + c = 0 ... rest is not difficult.

OpenStudy (anonymous):

ok so frm quadratic method i shall find out the value of p?

OpenStudy (experimentx):

yes ... use the variable separation after that.

OpenStudy (experimentx):

you will give two values of 'p' ... and it will give you two different solutions.

OpenStudy (anonymous):

ok i got it thnk u so much! :)

OpenStudy (experimentx):

yw